### v1

# TIP-Search: Time-Predictable Inference Scheduling for Market Prediction under Uncertain Load

Abstract
We present TIP-Search, a deadline-aware inference scheduling framework for real-time market prediction under uncertainty. In latency-critical environments such as high-frequency trading, inference decisions must be made within strict time budgets despite variable model performance and heterogeneous data distributions. TIP-Search dynamically selects one model per task from a pool of pre-trained deep learning models, balancing inference latency and predictive accuracy. Our approach leverages offline profiling and domain-aware generalization to route each task to the fastest eligible model expected to meet its deadline. Extensive evaluations on three real-world limit order book (LOB) datasets—FI-2010, Binance BTC/USDT, and LOBSTER AAPL—demonstrate that TIP-Search achieves up to 8.5% higher accuracy than fixed or random strategies while satisfying 100% of deadline constraints. We further conduct comprehensive ablation studies and analyze latency–throughput tradeoffs across scheduling policies.
Index Terms:
Real-time inference, deadline scheduling, heterogeneous models, market prediction, accuracy–latency tradeoff, LOB datasets
I Introduction
Modern financial systems increasingly rely on deep learning models for tasks such as limit order book (LOB) prediction, algorithmic trading, and price movement forecasting. These applications often operate under stringent latency constraints, where decision quality is tightly coupled with response time. The accuracy–latency tradeoff therefore becomes a central consideration in real-time deployment.
Deep learning inference, however, is inherently variable in latency, especially when models differ in size, architecture, and domain specialization. For instance, a convolutional model trained on high-frequency data from Binance may offer low-latency predictions but perform poorly when evaluated on Nasdaq-based FI-2010 data. Conversely, larger or more robust models may generalize well across market regimes yet incur higher computational cost. In high-stakes scenarios where each millisecond can alter trading outcomes, the need to reconcile inference accuracy with temporal predictability becomes imperative.
Traditional approaches to model deployment in financial AI pipelines typically assume fixed model allocation, offline provisioning for worst-case latency, or simplified selection heuristics. These methods fail to adapt to real-time variability in both data characteristics and system performance. Market environments are non-stationary, exhibiting bursty task arrivals and shifting distributions. In such settings, static selection strategies often lead to suboptimal performance: either violating latency guarantees or compromising predictive accuracy due to model–data mismatch.
This work addresses the problem of online model selection under deadline constraints by introducing TIP-Search, a framework for Time-Predictable Inference Scheduling. Rather than assigning a single model to all incoming tasks, TIP-Search dynamically selects a model per task from a heterogeneous pool, each trained on a distinct real-world financial dataset. The selection process is guided by offline latency profiling and estimated generalization performance, enabling per-task routing to the fastest eligible model with the highest expected accuracy.
TIP-Search is built on the insight that the inference-time–accuracy landscape forms a Pareto frontier. When sufficient slack is available, the scheduler favors more accurate models; under tighter deadlines, it prioritizes those with faster response time. Crucially, this tradeoff is navigated without violating task-level timing constraints.
LOB Task InputDeadline MonitorModel SelectorModel A(Low Latency)Model B(Balanced)Model C(High Accuracy)Prediction OutputModel Pool
Figure 1: TIP-Search architecture: Real-time LOB tasks are dispatched to a model pool based on deadline monitoring and selector logic.
To evaluate the efficacy of TIP-Search, we conduct experiments using three publicly available financial datasets: FI-2010, Binance BTC/USDT, and LOBSTER AAPL. Each dataset is used to train one DeepLOB model, forming a heterogeneous ensemble. Using realistic task arrival rates and fixed deadlines, we show that TIP-Search consistently improves prediction accuracy over fixed-model and random-selection baselines while maintaining 100% deadline satisfaction.
This paper makes the following contributions. First, we formulate the problem of deadline-constrained model selection in the context of real-time market prediction. Second, we propose TIP-Search, a practical and modular scheduling framework that combines model latency profiling with source-aware accuracy estimation. Third, we demonstrate through comprehensive evaluation that TIP-Search outperforms conventional heuristics across multiple datasets and decision conditions. Finally, we provide ablation analyses and design insights to support deployment in high-frequency inference environments.
To the best of our knowledge, TIP-Search represents the first attempt to jointly optimize for inference accuracy and real-time deadline satisfaction in the context of financial deep learning under uncertainty. The proposed framework contributes to a growing body of work on robust AI scheduling, and opens new directions for time-aware inference orchestration in latency-critical domains.
II Background and Motivation
Real-time financial decision systems, such as high-frequency trading and limit order book (LOB) prediction engines, operate under stringent temporal constraints. In such settings, even minor delays in model inference can lead to missed market opportunities or suboptimal trades. Consequently, the deployment of deep learning models in these environments must address not only the predictive performance of each model but also its computational latency. The core operational requirement is the ability to deliver high-quality predictions within strict and deterministic timing bounds.
Despite significant advances in model design for financial time series prediction, the integration of these models into real-time systems remains challenging. Inference latency varies considerably across different architectures and deployment configurations due to differences in computational complexity, parameter size, and hardware contention. Additionally, task input characteristics—such as the volatility of market dynamics or the structural patterns of LOB events—can influence inference behavior, leading to heterogeneous response times even within the same model.
Crucially, financial inputs arrive in an inherently stochastic manner. Unlike domains with stable periodic input rates, market data streams are characterized by burstiness and temporal irregularity, resulting from shifts in trader behavior, economic announcements, or liquidity shocks. This temporal uncertainty amplifies the difficulty of static resource allocation or fixed pipeline designs. Furthermore, the assumption that a single model can generalize across all market conditions and asset classes is often violated in practice. For instance, a model trained on historical Nasdaq data may perform poorly when evaluated on cryptocurrency order books, not due to architectural flaws but due to domain shift.
Previous attempts to mitigate these issues have largely focused on two types of strategies. The first class adopts conservative worst-case latency provisioning, in which all tasks are routed through the fastest model to guarantee deadline compliance. While this approach simplifies scheduling, it typically sacrifices prediction quality, especially when faster models are less expressive or trained on less relevant data. The second class relies on offline model selection, where a fixed model is assigned to all tasks under the assumption of input homogeneity. Such approaches ignore runtime variability and fail to adapt to changes in input source or model availability.
This paper addresses these limitations by proposing a dynamic and adaptive framework—TIP-Search—that selects the optimal model for each task at runtime based on latency constraints and expected predictive accuracy. The motivation is rooted in the observation that different models offer distinct tradeoffs in the latency–accuracy space, and that effective real-time inference demands a scheduling policy capable of navigating this tradeoff dynamically. TIP-Search avoids both the pessimism of worst-case provisioning and the rigidity of static model assignment by using a lightweight decision mechanism to select the fastest eligible model that is likely to yield the most accurate result, given the task’s characteristics and timing budget.
In summary, the motivation for TIP-Search arises from the confluence of three key challenges in real-time market inference: variability in model latency, heterogeneity in model generalization, and uncertainty in input arrival patterns. These challenges necessitate a fundamentally different approach to model scheduling—one that explicitly accounts for both performance and timing constraints while remaining adaptable to runtime conditions. The TIP-Search framework is designed to meet this need, providing a practical and theoretically grounded solution for real-world financial inference systems.
III Problem Formulation
Notation
To support formal analysis, we define key symbols in Table I.
TABLE I: Key Symbols and Definitions
Symbol Description
$t_{i}$ Input task (LOB snapshot) arriving at time $a_{i}$
$a_{i}$ Arrival time of task $t_{i}$
$d_{i}$ Deadline for task $t_{i}$, defined as $a_{i}+\Delta$
$\Delta$ Maximum allowable inference latency (e.g., 10ms)
$\mathcal{M}$ Set of all available models, $\{M_{1},M_{2},\dots,M_{K}\}$
$M_{k}$ $k$-th model in the pool $\mathcal{M}$
$L_{k}$ Inference latency of model $M_{k}$
$A_{k}(x)$ True (unknown) accuracy of $M_{k}$ on input $x$
$\hat{A}_{k}(t_{i})$ Estimated accuracy of $M_{k}$ for task $t_{i}$
$\mathcal{F}(t_{i})$ Models satisfying $a_{i}+L_{k}\leq d_{i}$
$M^{\star}(t_{i})$ Selected model maximizing $\hat{A}_{k}(t_{i})$
Problem Setting
We consider a real-time inference scenario in which a financial prediction engine receives a continuous stream of input tasks, each corresponding to a limit order book (LOB) snapshot. Each task $t_{i}$ arrives at time $a_{i}$ and must be processed before its deadline $d_{i}=a_{i}+\Delta$, where $\Delta$ is a fixed system-wide latency budget.
The system is equipped with a model pool =$\mathcal{M}=\{M_{1},M_{2},\ldots,M_{K}\}$, where each $M_{k}$ is trained on different financial datasets or under varying architectures. Each model is associated with an offline-measured latency $L_{k}$ and an (unknown at runtime) input-conditional accuracy $A_{k}(x)$. To facilitate scheduling, $\hat{A}_{k}(t_{i})$ denotes an estimated accuracy, obtained via domain heuristics or validation feedback.
A key challenge arises from the trade-off between inference speed and accuracy. Faster models may underperform on out-of-distribution inputs, while more accurate models may exceed per-task latency budgets. Additionally, task inputs do not include labels or metadata indicating which model is optimal, which compounds the uncertainty in selection.
Optimization Objective
The objective is to maximize the total number of correct predictions across all input tasks while satisfying their timing constraints. Formally, the optimization problem is expressed as:
$\max\quad$ $\sum_{i}\mathbb{I}[\text{correct}(t_{i},M_{k})]$ (1)
s.t. $a_{i}+L_{k}\leq d_{i},\quad\forall i$ (2)
where $\mathbb{I}[\cdot]$ denotes the indicator function returning 1 if the prediction made by $M_{k}$ for $t_{i}$ is correct.
This formulation assumes that latency values $L_{k}$ are deterministic and accurately profiled. However, the true accuracy $A_{k}(x)$ remains unknown at decision time, motivating the use of accuracy estimators $\hat{A}_{k}(t_{i})$ derived from task features or past performance.
TIP-Search Policy Formulation
TIP-Search approaches this problem by first filtering out models that violate the deadline constraint, and then selecting the one with the highest expected accuracy. This decision process is formalized as:
$M^{\star}(t_{i})=\arg\max_{M_{k}\in\mathcal{F}(t_{i})}\hat{A}_{k}(t_{i}),$ (3)
where $\mathcal{F}(t_{i})=\{M_{k}\in\mathcal{M}\mid a_{i}+L_{k}\leq d_{i}\}$. By combining offline latency profiling with runtime accuracy estimation, TIP-Search enables per-task model selection under uncertainty, ensuring deadline compliance while pursuing optimal prediction quality.
IV Theoretical Analysis
This section provides a formal study of TIP-Search’s schedulability, computational complexity, and accuracy gap relative to an idealized oracle policy.
IV-A Schedulability Guarantee
We define a task $t_{i}$ as schedulable if there exists at least one model $M_{k}$ such that $L_{k}\leq\Delta$. Assuming model latencies are fixed and known offline, we obtain the following result:
Theorem 1 (Feasibility Guarantee).
If $\exists M_{k}\in\mathcal{M}$ such that $L_{k}\leq\Delta$, then every task $t_{i}$ will be assigned to a model that satisfies its deadline $d_{i}=a_{i}+\Delta$.
Proof.
TIP-Search constructs the eligible model set $\mathcal{E}(t_{i})=\{M_{k}\mid L_{k}\leq\Delta\}$. Since this set is non-empty by assumption, the algorithm selects the model with highest estimated accuracy $\hat{A}_{k}(t_{i})$ among the eligible models. Thus, the chosen model $M^{*}(t_{i})$ satisfies $L_{M^{*}(t_{i})}\leq\Delta$, guaranteeing $a_{i}+L_{M^{*}(t_{i})}\leq d_{i}$. ∎
IV-B Runtime Complexity
Let $K=|\mathcal{M}|$. TIP-Search performs:
  • $\mathcal{O}(K)$ operations to filter deadline-eligible models;
  • $\mathcal{O}(K)$ operations to compute $\arg\max\hat{A}_{k}(t_{i})$.
Hence, the total per-task runtime is $\mathcal{O}(K)$—constant for bounded-size model pools.
IV-C Accuracy Gap to Oracle Policy
Let $M^{\text{oracle}}(t_{i})$ denote the optimal selector that uses the true (but unknown) accuracy $A_{k}(t_{i})$:
Definition 1 (Oracle Selection Policy).
$M^{\text{oracle}}(t_{i})=\arg\max_{M_{k}\in\mathcal{E}(t_{i})}A_{k}(t_{i})$
TIP-Search instead uses:
$M^{*}(t_{i})=\arg\max_{M_{k}\in\mathcal{E}(t_{i})}\hat{A}_{k}(t_{i})$
Let cumulative regret over $n$ tasks be:
$\mathcal{R}_{n}=\sum_{i=1}^{n}\left(A_{M^{\text{oracle}}(t_{i})}(t_{i})-A_{M^{*}(t_{i})}(t_{i})\right)$
Theorem 2 (Bounded Regret under Estimation Error).
If $\lvert\hat{A}_{k}(t_{i})-A_{k}(t_{i})\rvert\leq\epsilon$ for all $k$ and $t_{i}$, then:
$\mathcal{R}_{n}\leq n\cdot\epsilon$
Proof.
At each task $t_{i}$, TIP-Search may miss the optimal model by at most $\epsilon$ in accuracy. Thus, the total regret across $n$ tasks is bounded by $n\cdot\epsilon$. ∎
V TIP-Search Algorithm Design
TIP-Search is designed as a deadline-aware inference orchestration framework, aimed at ensuring predictable response under uncertainty. It operates by dynamically selecting a model from a heterogeneous pool such that the execution satisfies latency constraints while maximizing expected predictive accuracy.
V-A System Architecture
The system processes real-time tasks, each corresponding to a market snapshot, and performs deadline-constrained inference using a multi-model dispatch mechanism. The architecture comprises three key components: (1) a deadline monitor that computes the residual slack for each task; (2) a model selector that filters and ranks eligible models according to latency constraints and estimated accuracy; and (3) a model pool consisting of multiple pre-trained models, each with distinct latency–accuracy tradeoffs.
Figure 1 illustrates the TIP-Search workflow. When a new task arrives, its deadline is computed based on the arrival time and a fixed budget. The scheduler evaluates all available models for latency eligibility and selects the one with the highest expected accuracy among them.
V-B Latency Profiling and Accuracy Estimation
Prior to deployment, all models are profiled to record their average latency under warm-up conditions. These latency profiles $L_{k}$ are stored for all models $M_{k}\in\mathcal{M}$ and referenced at runtime to enforce deadline compliance. For predictive quality, each model is also associated with an empirical accuracy estimator $\hat{A}_{k}(t_{i})$, which may be derived from offline validation or adaptive online tracking.
V-C Scheduling Logic
Given a task $t_{i}$ arriving at time $a_{i}$ and deadline $d_{i}$, the scheduler identifies the set $\mathcal{E}$ of models for which $a_{i}+L_{k}\leq d_{i}$ holds. Among the eligible models, it selects $M^{*}(t_{i})$ that maximizes $\hat{A}_{k}(t_{i})$. The scheduling strategy ensures that all dispatched tasks respect timing constraints.
Algorithm 1 TIP-Search: Deadline-Aware Scheduling Policy
1:Task $t_{i}$ with arrival $a_{i}$ and deadline $d_{i}$
2:$\mathcal{F}(t_{i})\leftarrow\emptyset$
3:for each model $M_{k}\in\mathcal{M}$ do
4:     if $a_{i}+L_{k}\leq d_{i}$ then
5:         $\mathcal{F}(t_{i})\leftarrow\mathcal{F}(t_{i})\cup\{M_{k}\}$
6:     end if
7:end for
8:if $\mathcal{F}(t_{i})\neq\emptyset$ then
9:     $M^{\star}(t_{i})\leftarrow\arg\max_{M_{k}\in\mathcal{F}(t_{i})}\hat{A}_{k}(t_{i})$
10:     Dispatch $t_{i}$ to $M^{\star}(t_{i})$
11:else
12:     Drop $t_{i}$ or invoke fallback
13:end if
V-D Execution Workflow
At runtime, the inference loop continuously monitors incoming tasks. Upon each arrival, the system computes the remaining time to deadline and invokes the scheduling logic to determine the model to use. The selected model processes the task immediately. Since all latency profiles are preloaded and accuracy estimators are lightweight, the scheduling overhead is negligible.
V-E Integration and Extension
The modularity of TIP-Search allows it to be integrated into existing inference engines. It is agnostic to model architecture and supports GPU/CPU inference interchangeably. For extensibility, the accuracy estimation component can incorporate confidence scores, domain embeddings, or self-supervised feedback, making it adaptable to future research in learning-based scheduling.
V-F Theoretical Analysis and Feasibility Considerations
To analyze the schedulability of TIP-Search under hard real-time constraints, we define the notion of task feasibility with respect to the system’s model pool $\mathcal{M}$. Given a task $t_{i}$ with arrival time $a_{i}$ and deadline $d_{i}$, TIP-Search guarantees feasibility if there exists at least one model $M_{k}$ such that $a_{i}+L_{k}\leq d_{i}$. This ensures that task-level schedulability is preserved provided the model latency profiles $\{L_{k}\}$ are known and stable.
In the worst case, the selection logic incurs $\mathcal{O}(K)$ computational overhead per task, where $K$ is the number of candidate models. This ensures constant-time decision latency for bounded $K$, which holds in typical real-time deployments with pre-enumerated inference configurations.
Moreover, we observe that the scheduling decision reduces to a constrained maximization over a pre-filtered set, which can be formulated as a special case of deadline-constrained utility maximization. While the exact optimality of the surrogate accuracy function $\hat{A}_{k}$ cannot be guaranteed in non-stationary settings, TIP-Search is provably deadline-feasible given static latency bounds.
VI Formal System Guarantees
In this section, we provide formal guarantees on the schedulability and responsiveness of TIP-Search under uncertain and bursty load conditions. We also analyze the interference robustness of TIP-Search when multiple models are co-located on shared inference hardware.
VI-A Schedulability under Bursty Arrivals
Let $\lambda_{\text{max}}$ denote the maximum arrival rate (tasks per second), and $\mathbf{L_{\min}}=\min_{k}L_{k}$ denote the minimum model latency. For a system with strict deadline $\Delta$ and $K$ pre-profiled models, TIP-Search ensures feasibility under the following condition:
Theorem 3 (Schedulability under Peak Load).
If $\lambda_{\text{max}}\cdot\mathbf{L_{\min}}\leq 1$, then TIP-Search guarantees 100% deadline satisfaction under bursty arrivals.
Proof.
Since TIP-Search always selects the fastest eligible model $M_{k}$ satisfying $L_{k}\leq\Delta$, the system can serve up to $1/L_{k}$ tasks per second per model. Under worst-case load, if $\lambda_{\text{max}}\leq 1/L_{\min}$, at least one model remains schedulable at all times. ∎
VI-B Response-Time Upper Bound
Let $R_{i}$ denote the end-to-end response time for task $t_{i}$ under TIP-Search scheduling. Then:
Theorem 4 (Response Time Bound).
Assuming zero queuing delay and deterministic latency $L_{k}$ per model, we have:
$R_{i}\leq\min_{M_{k}\in\mathcal{F}(t_{i})}L_{k}$
Proof.
By construction, TIP-Search filters models satisfying $a_{i}+L_{k}\leq d_{i}$ and chooses the one with highest $\hat{A}_{k}(t_{i})$. Thus, $L_{k}\leq\Delta$ and $R_{i}\leq L_{k}$ for the selected model. ∎
VI-C Robustness to Co-located Interference
In multi-model deployments, concurrent tasks may share inference hardware, leading to performance interference. TIP-Search is robust to such contention due to its latency-awareness. We formalize this as:
Definition 2 (Contention-Aware Schedulability).
Let $\tilde{L}_{k}=L_{k}+\delta_{k}$ be the observed latency under co-located interference. If $\exists M_{k}$ such that $\tilde{L}_{k}\leq\Delta$, TIP-Search preserves deadline feasibility.
Theorem 5 (Interference-Robust Scheduling).
If the interference margin $\delta_{k}$ satisfies $\max_{k}\delta_{k}\leq\epsilon$ and $\min_{k}L_{k}+\epsilon\leq\Delta$, then TIP-Search remains deadline-feasible.
Proof.
Follows directly from the eligibility check in the TIP-Search selection policy, which can be extended to use $\tilde{L}_{k}$ at runtime. ∎
VI-D Robustness under DVFS and Energy Constraints
Modern embedded systems often use dynamic voltage and frequency scaling (DVFS) to reduce power consumption. Let $L_{k}(f)$ denote the latency of model $M_{k}$ at frequency $f$. Under scaling, TIP-Search remains feasible if:
Theorem 6 (DVFS Compatibility).
If there exists frequency $f$ such that $\min_{k}L_{k}(f)\leq\Delta$, then TIP-Search preserves feasibility under DVFS.
Proof.
Offline profiling must include $L_{k}(f)$ at various $f$. During runtime, TIP-Search can select the optimal $(M_{k},f)$ pair such that $L_{k}(f)\leq\Delta$, while maximizing $\hat{A}_{k}(t_{i})$. ∎
VI-E Resource Isolation in Multi-GPU Scheduling
In multi-GPU systems, different models may be bound to disjoint GPU cores. Let $G_{j}$ denote GPU $j$, and $\mathcal{M}_{j}$ the model set resident on $G_{j}$. Then:
Theorem 7 (GPU Isolation Guarantee).
If each GPU $G_{j}$ hosts models with $L_{k}\leq\Delta$, and task routing ensures $t_{i}\mapsto G_{j}$ iff $M_{k}\in\mathcal{M}_{j}$ is feasible, TIP-Search guarantees per-GPU deadline compliance.
Proof.
Per-GPU feasibility follows from intra-GPU eligibility checks and independence of scheduling across $G_{j}$. ∎
VII Experimental Evaluation
This section presents an empirical evaluation of TIP-Search under deadline-constrained market prediction workloads. The experiments are designed to assess the framework’s ability to maintain real-time responsiveness while maximizing predictive accuracy, as well as to quantify its generalization capacity, scheduling efficiency, and comparative performance against representative baselines.
VII-A Experimental Setup and Datasets
To simulate realistic online inference scenarios, we construct a heterogeneous workload using three distinct financial limit order book (LOB) datasets. The FI-2010 dataset [3] comprises ten days of Nasdaq LOB data annotated with five-class mid-price movements. The Binance BTC/USDT dataset contains a 2025/05/10 snapshot of raw cryptocurrency trade events labeled for three-class price direction. The LOBSTER AAPL dataset provides high-frequency order book data for Apple Inc. over ten trading days, representing a more stable equity market.
Each dataset is used to independently train a DeepLOB model [4], yielding three heterogeneous models in the inference pool. These models are deployed on a single GPU, and runtime inference is triggered every 5ms, with a strict 10ms deadline per task. Each task includes an LOB snapshot, a ground-truth directional label, and an identifier of the dataset source. The TIP-Search framework must select a model per task to maximize predictive performance without violating the deadline constraint.
VII-B Evaluation Protocol and Metrics
We evaluate all scheduling policies on a shared stream of 298 tasks, sampled evenly across the three datasets. Each policy operates in isolation under identical inter-arrival and hardware conditions. Four primary metrics are collected: prediction accuracy under deadline constraint, deadline satisfaction rate, average inference latency per model, and total task throughput. These jointly reflect both effectiveness and efficiency of the tested strategies.
VII-C Generalization Performance Across Domains
The generalization behavior of individual models is examined by constructing an accuracy matrix where each row corresponds to the input source domain and each column to a deployed model. The results, visualized in Figure 2, reveal significant asymmetries. The Binance and LOBSTER models exhibit strong cross-domain accuracy, while the FI-2010 model demonstrates poor generalization when applied to unseen market distributions. These discrepancies underscore the importance of dynamic model selection in heterogeneous environments.
Figure 2: Cross-domain generalization matrix: accuracy of each model across source domains.
TIP-Search leverages these asymmetries by adaptively routing tasks to models with higher expected generalization accuracy under runtime constraints. The observed routing patterns confirm that model selection in TIP-Search aligns with latent data characteristics even without access to explicit source labels during execution.
VII-D Strategy-Level Comparative Analysis
To quantify the benefit of adaptive scheduling, we compare TIP-Search against three baseline policies: a fixed-model strategy that always dispatches to one model, a random selection strategy among deadline-eligible models, and a source-aware strategy that assumes oracle access to ground-truth dataset labels. Figure 4 presents the overall accuracy of each strategy across the three domains.
VII-E Comparison with Classical Real-Time Scheduling
To establish a connection with classical real-time scheduling theory, we simulate two canonical policies: Earliest Deadline First (EDF) and Least Laxity First (LLF). Although these strategies are traditionally used for CPU-level task dispatching rather than inference model selection, they provide valuable baselines for interpreting system responsiveness under dynamic load.
In our adaptation, EDF selects the model with the shortest remaining slack (i.e., $d_{i}-a_{i}-L_{k}$), while LLF uses the laxity ratio to rank candidate models. Both strategies are evaluated on the same 298-task stream. The resulting accuracy under deadline constraints was inferior to TIP-Search, suggesting that incorporating domain-specific accuracy estimation provides meaningful gains over purely temporal heuristics.
Figure 3: Comparison of TIP-Search with classical EDF and LLF strategies under fixed deadline constraints.
Figure 4: Unified strategy-domain heatmap: accuracy of each dispatch policy across source domains.
TIP-Search consistently achieves higher accuracy than all three baselines. While the fixed-model and random strategies suffer from mismatches between task and model domain, the source-aware baseline performs well only when the optimal model aligns exactly with the source, which is not always the case. TIP-Search surpasses even this oracle-guided strategy by balancing accuracy against timing constraints in real time.
VII-F Latency and Throughput Under Real-Time Constraints
To assess the efficiency of scheduling policies, we analyze the end-to-end inference latency and system throughput for each model. Figure 5 shows that the Binance model offers the lowest latency, while the FI-2010 model incurs significantly higher computational cost. TIP-Search adaptively avoids high-latency models when deadlines are tight, resulting in low variance in end-to-end inference time and stable throughput across bursts.
Figure 5: Latency and throughput profile of each model under controlled real-time workloads.
Despite its dynamic nature, TIP-Search achieves an average throughput exceeding 190 tasks per second while preserving 100
VII-G Ablation Study: Fixed, Random, and Source-Only Policies
To isolate the contribution of adaptive accuracy estimation, we conduct an ablation study comparing three simplified variants. Figures 68 present per-domain heatmaps showing the task-model assignment accuracy under each baseline.
Figure 6: Fixed strategy: accuracy heatmap under static model selection.
Figure 7: Random strategy: accuracy heatmap with uniform dispatch among eligible models.
Figure 8: Source-only strategy: task assignment based on known origin domain.
Across all domains, TIP-Search outperforms the baselines by avoiding misaligned task-model pairs. The fixed strategy yields the lowest accuracy due to model bias. The random strategy performs slightly better but fails to exploit structure in task arrival. The source-only strategy offers competitive performance when task domain matches training data, but degrades when domain alignment is insufficient. TIP-Search maintains high accuracy under all settings by jointly reasoning about deadline feasibility and generalization profiles.
VII-H Summary and Observations
The experimental results validate TIP-Search as a robust framework for deadline-aware scheduling under uncertainty. It achieves high prediction accuracy without deadline violations, generalizes across domains in the absence of source labels, and introduces minimal scheduling latency. These characteristics make it a compelling candidate for real-time inference in finance, robotics, and other high-frequency decision systems.
VIII Ablation Study and Discussion
To assess the effectiveness of TIP-Search in real-time inference, we conduct ablation studies that compare its performance with simplified decision strategies. Each strategy imposes a constraint on the model selection logic, thereby isolating the impact of specific scheduling principles.
VIII-A Experimental Conditions
The ablation experiments are structured to examine four distinct policies under identical workload conditions. The Fixed strategy consistently employs a single preselected model for all tasks, regardless of input characteristics. The Random strategy selects among eligible models with uniform probability, without consideration of model quality. The Source-Only strategy dispatches each task to the model trained on the same source as the input data. Finally, TIP-Search—our proposed framework—dynamically selects the model with the highest expected accuracy from the subset of models that satisfy the deadline constraint.
Each policy is evaluated on a unified stream of market tasks drawn from multiple domains. Every task must be executed within a 10ms timing budget, and all experiments are executed under identical hardware and scheduling conditions to ensure comparability.
VIII-B Quantitative Results
Table II presents the empirical outcomes in terms of predictive accuracy and deadline satisfaction across all strategies. The results indicate that TIP-Search consistently achieves the highest accuracy (51.01%) while preserving 100% deadline compliance. In contrast, fixed models suffer from domain-specific limitations. For example, although the Lobster-trained model achieves 57.05% accuracy, it fails to generalize across sources. The FI-2010 model, by comparison, shows markedly poor performance (10.40%), highlighting the pitfalls of single-source deployment. The random strategy yields moderate accuracy (45.64%) due to probabilistic exposure to both good and bad models, while source-only routing underperforms due to its inability to recognize generalization failure.
TABLE II: Performance Comparison of Scheduling Strategies
Strategy Accuracy (%) Deadline Met (%)
Fixed (Binance) 43.96 100.0
Fixed (FI-2010) 10.40 100.0
Fixed (Lobster) 57.05 100.0
Random Selection 45.64 100.0
Source-Based Routing 39.26 100.0
TIP-Search (Ours) 51.01 100.0
VIII-C Visualization of Scheduling Efficacy
To further elucidate the effectiveness of each strategy, Figure 9 presents a heatmap of task source versus selected model accuracy. It illustrates that TIP-Search adapts to input variability by selecting high-performing models per task, thereby avoiding low-accuracy assignments such as those frequently triggered by the FI-2010 model. In contrast, other strategies exhibit rigid or naive selection behaviors, leading to degraded accuracy on mismatched inputs.
Figure 9: Accuracy heatmap showing model-task pairings across scheduling strategies.
VIII-D Interpretation and Implications
These ablation results underscore several key findings. First, model mismatch remains a fundamental challenge: no single model performs consistently well across all domains. Second, static strategies—including fixed or source-specific dispatch—fail to account for generalization variability and thus underutilize the potential of heterogeneous model pools. Finally, TIP-Search demonstrates superior adaptability by leveraging deadline-aware inference and accuracy estimation to navigate tradeoffs in real time. This robustness is critical for deployment in latency-sensitive applications such as algorithmic trading and financial forecasting.
The subsequent section discusses broader implications for system-level integration and outlines potential extensions of TIP-Search to multi-GPU, edge, or federated environments.
IX Deployment Considerations and Limitations
IX-A Real-Time System Integration
TIP-Search is architected to operate with minimal runtime overhead, making it amenable to integration with existing real-time financial inference pipelines. The scheduling mechanism executes entirely on the CPU, thereby avoiding contention with GPU-based inference workloads and preserving parallelism. In deployment environments with multiple resident models, TIP-Search supports concurrent GPU model sharing by leveraging preloaded weights in memory, which significantly reduces switching and loading latencies. Furthermore, the system assumes a warm-start initialization for all models, ensuring that cold-start latency penalties are eliminated prior to runtime. In dynamic market conditions, TIP-Search remains deadline-aware and supports per-task timing adaptation, enabling seamless operation under varying load intensities or application-specific latency constraints.
IX-B Scalability in Heterogeneous Environments
The proposed scheduling framework exhibits favorable scalability characteristics when managing a moderate number of candidate models, typically ranging from ten to twenty. This scale suffices for most real-time inference scenarios in market prediction, where models are often trained on distinct market instruments, time frames, or architectural variants. Nevertheless, scaling beyond this range may introduce computational overhead during eligibility filtering and accuracy ranking. In such cases, performance can be preserved through hierarchical scheduling mechanisms, lightweight surrogate scoring models, or pruning strategies that narrow the candidate set based on early-exit conditions or hardware-specific latency predictors.
IX-C Limitations and Future Directions
Despite its effectiveness in balancing inference quality and latency, TIP-Search is subject to several limitations. First, the system relies on the availability of multiple pre-trained models, each tailored to a specific domain or data distribution. This requirement introduces additional training and maintenance cost, particularly in evolving market environments. Second, TIP-Search currently assumes stationarity in model generalization profiles over time; however, financial markets are inherently non-stationary. Without mechanisms for continual learning or model retraining, TIP-Search may experience degradation in performance due to regime shifts or concept drift. Finally, the system estimates model accuracy using static or heuristically updated metrics, which may be insufficient under adversarial conditions or in the presence of previously unseen input distributions.
To address these challenges, future extensions of TIP-Search could incorporate online adaptation mechanisms, including self-supervised performance feedback, domain shift detection, or ensemble uncertainty quantification. Additionally, integrating low-cost latency prediction models or confidence-aware inference strategies may further enhance the system’s responsiveness in highly dynamic environments.
IX-D Transferring to Embedded Inference Scenarios
Although this work focuses on financial inference systems, the TIP-Search framework is well-suited for deployment in embedded real-time environments. In embedded AI systems, such as autonomous drones, mobile robots, and sensor-driven industrial controllers, real-time constraints and heterogeneous inference paths are common challenges.
TIP-Search’s model-pool-based scheduling policy generalizes to scenarios where different model configurations correspond to distinct memory footprints, energy consumption profiles, or data preprocessing pipelines. For example, a robotic system may maintain one small CNN for low-latency edge detection and another transformer-based model for more accurate, but slower, spatial reasoning. TIP-Search can dynamically choose the inference path based on a task’s urgency and timing budget.
Moreover, the offline latency profiling strategy aligns with the static analysis common in embedded systems, while the lightweight runtime dispatch ensures minimal CPU overhead. Because embedded systems often operate under constrained computational budgets, TIP-Search’s $\mathcal{O}(K)$ decision complexity and deadline-satisfaction guarantee offer practical benefits.
In future extensions, TIP-Search could integrate with existing real-time operating systems (RTOS) via inference-aware task schedulers, or be combined with early-exit networks to support dynamic voltage and frequency scaling (DVFS) in power-sensitive deployments.
X Relevance to Real-Time Systems Research
Although TIP-Search is evaluated in the context of financial market prediction, the methodology and scheduling challenges it addresses are fundamental to the broader class of real-time AI systems. The proposed framework deals with online task arrival, strict latency constraints, heterogeneous execution paths, and dynamic model selection—elements commonly encountered in robotics, autonomous vehicles, embedded perception, and other cyber-physical applications.
In particular, the formulation of latency-aware, accuracy-guided model dispatch closely mirrors decision-making under bounded computational budgets in real-time embedded systems. Moreover, TIP-Search aligns with the traditional goals of real-time scheduling: deadline satisfaction, throughput maximization, and resource-aware dispatch.
By framing the deep learning inference pipeline as a real-time scheduling problem with accuracy–latency trade-offs, this work extends the RTSS discourse into the AI-inference domain. It demonstrates that rigorous real-time techniques are not only applicable but also beneficial in managing uncertainty and dynamics in data-driven settings. We believe TIP-Search opens new directions for integrating RT scheduling theory with adaptive model-based decision systems.
XI Conclusion
We proposed TIP-Search, a time-predictable inference scheduling framework that adaptively selects deep models for real-time financial market prediction. TIP-Search enables robust accuracy under deadline constraints by leveraging domain-aware generalization and latency profiling.
Through experiments on three real-world datasets (Binance, FI-2010, LOBSTER), we showed that TIP-Search outperforms fixed, random, and source-specific strategies. Our ablation studies confirm its advantage in handling model diversity and latency variation.
TIP-Search provides a practical and modular enhancement to real-time AI systems, especially in latency-sensitive domains such as trading, robotics, and cyber-physical systems.